Write the null and alternative hypothesis of given problem:


    1. It is claimed that the average weight of a bag of biscuit is 250 grams with a standard deviation of 20.5 grams. Would you agree to this claim if a random sample of 50 bags of biscuits showed an average weight of 240 grams, using a 5% level of significance.



1
Expert's answer
2021-09-29T00:09:52-0400

Solution:

Null hypothesis, "H_0:\mu=250"

Alternative hypothesis, "H_1:\mu\ne250"

Assume that the population standard deviation is "\sigma=20.5" . The population standard deviation is known and the sample size n is large "(n \geq 30)" . Hence, the test statistic is "z=\frac{\bar{x}-\mu}{\sigma / \sqrt{n}}" , and the value is "z=\frac{\bar{x}-\mu}{\sigma / \sqrt{n}}=\frac{240-255}{20.5 / \sqrt{50}}=-5.174" . The test is two tailed (because H1 contains inequality "\neq" ) and the critical values are "z_{-0.025}=-1.96" and "z_{0.025}=1.96" . Since "z<-z_{0.025}" , i.e. the z lies to the left of the critical value, we reject the null hypothesis.




We would not agree with a claim, the true average weight of a bag of biscuits is not 250 grams. Answer: the true average weight of a bag of biscuits is not 250 grams.